Be sure to answer all parts. The pH of a 0.30 M solution of a weak base is 10.02. What is the Kb of the base? Kb = × 10
Answer #1

Let us consider weak base be BOH

BOH(aq) -------> B+(aq) + OH-(aq)

pH + pOH = 14

pOH = 14-10.02 = 3.98

pOH = -log [OH-] = 3.98

[OH-] = 10-3.98 = 0.000105 M

   BOH(aq) -------> B+(aq) +    OH-(aq)

Initial Concentration :    0.3 0 0

Change in concentration :(0.3-0.000105) +0.000105 +0.000105

Equilibrium concentration : 0.299895    0.000105    0.000105

Kb = [OH-] [B+]/[BOH] = (0.000105)2/0.299895 = 0.000000011/0.299895 = 3.67 X 10-8